x=-2 and x=0
-3x^3+7x^2-3x-2

Answer by checkley75(3666)

-3X^3+7X^2-3X-2 WHEN X=-2 THEN
-3(-2)^3+7(-2)^2-3(-2)-2
-3(-8)+7(4)+6-2
24+28+6-2=56 ANSWER.
------------------------------------------
WHEN X=0 THEN
-3(O)^3+7(0)^2-3(0)-2
-3*0+7(0)-0-2
-0+0-0-2=-2 ANSWER.